OpenStudy (anonymous):

Please Can You help? In Sardinia, approximately one in every 200 people has β-thalassemia. Using the mathematics from lecture, estimate the following: • a)The frequency of the β-thalassemia allele in the Sardinian population. Please express your answer as a number between 0 and 1 using one significant digit only. • b)The frequency of β-thalassemia carriers (genotype Aa) in the Sardinian population. Please express your answer as a number between 0 and 1 using one significant digit only.

OpenStudy (anonymous):

a) let the probability of the allele be p then p*p =1/200 p = 0.07 b)probability of carrier = 2*0.07*(1-0.07) (because any of the two allele can have it) =0.1302 = 0.1 (approx)

OpenStudy (anonymous):

c) req. probability = prob of being a carrier * prob of having the allele in child = 0.1302*1/2 =0.0651 = 0.07(approx)